OpenStudy (blake57roger):

What value of x makes the equation true? –7.12 = –4.8 + x A. –11.92 B. –11.2 C. –2.32 D. 2.32

OpenStudy (blake57roger):

@amistre64

OpenStudy (amistre64):

any ideas on what to do?

OpenStudy (blake57roger):

a few

OpenStudy (amistre64):

let me here them

OpenStudy (blake57roger):

you add them somehow to get your answer

OpenStudy (amistre64):

youve got the right idea, lets make it more sound tho if we have 2 things that are the same, they are equal .... then we want to keep them equal. whatever we do to one side we have to do to the other. on one side we have -4.8 + x, but we only want the x part .... if we add 4.8 to it will that get rid of the extra numbers?

OpenStudy (amistre64):

-7.12 = -4.8 + x +4.8 +4.8 ----------------- ?? = 0 + x

OpenStudy (amistre64):

if we estimate, then the solution is about -7+5